OpenStudy (anonymous):

What is the number of real solutions? –11x2 = x + 11 cannot be determined one solution two solutions no real solutions

OpenStudy (anonymous):

I think it is cannot be determined... but I want someone to check my answer.

OpenStudy (anonymous):

rewrite: \[11x^2+x+11=0\] a = 11, b = 1, c = 11 \[b^2-4ac= 1-4(11)(11) <0 \Rightarrow \text{complex zeros}\]

OpenStudy (anonymous):

So there are zero solutions?

OpenStudy (anonymous):

no. have a look... http://www.mathwarehouse.com/quadratic/discriminant-in-quadratic-equation.php

OpenStudy (anonymous):

There are two real solutions!

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Negative Discriminant b² − 4ac < 0 No Real Solutions Two Imaginary Solutions

OpenStudy (anonymous):

there you go!

OpenStudy (anonymous):

So there are no real solutions?

OpenStudy (anonymous):

yep